S unit 5.2 MC 2A 2 01 000s of units) Refer to the dus tan abova. suppoirs ths hrn bs curenty prod ucits 200 ts of pioduct note that qaantity suppled ls measured In thousand s of an.The total ocast svins fesuking fron the lechnolos eal hange bed uial tos unclude a do llat ssn tn tront of yout anpwer, and do nor Inser aconmsor kave anyspaces In your answer for ceample, your anser systhousand dollars, show th 7300011 thete ate no gents do not Include Zelos te do notshow the as S 00000
Answer #1

Solution:

Notice that with technological change (or technological advancement since the cost has decreased, hence increasing efficiency), the marginal cost curve has shifted downwards. To find the total cost saved, we first find the marginal cost functions: the MC curves are straight lines, and using two points formula, we can find the required functions (denoting 1000s of units by x and costs by y)

Finding equation for MC0: straight line passing through (0, 4) and (2, 5.2)

Slope = (5.2 - 4)/(2 - 0) = 0.6

Equation: y - 4 = 0.6(x - 0)

y = 0.6x + 4

Or MC0 = 0.6Q + 4

Then, initial total cost, TC0 = integration of MC0 with respect to Q

TC0 = 0.6*Q2/2 + 4Q + c

With Q = 2000 units, TC0 = 0.6*(2000)2/2 + 4*2000 + c = $1,208,000+c

Similarly, finding equation of MC1: straight line passing through (0, 1) and (2, 1.2)

Slope = (1.2 - 1)/(2 - 0) = 0.1

Equation: y - 1 = 0.1*(x - 0) that is y = 0.1x + 1

Of MC1 = 0.1Q + 1

So, new total cost, TC1 = integration of MC1 with respect to Q

TC1 = 0.1*Q2/2 + Q + c

With Q = 2000 units, TC1= 0.1*(2000)2/2 + 2000 + c = $202,000+c

So, with fixed cost of $c same in both cases, total cost reduced (or saved) = TC0 - TC1

= 1208000 - 202000 = $1,006,000 (to be answered as $1006000)